# Zeroth-order log-concave sampling

### v1

@@ 1.1.2 Order-preserving warm-start generation via annealing
Perhaps the most relevant mathematical notion to tackle this question is Gross’ hypercontractivity of semigroups [Gro75] (see Proposition 3.1): for a semigroup $(P_{t})_{t\geq 0}$ reversible with respect to an invariant measure $\pi$, it is known that ($\mathsf{LSI}$) for $\pi$ is equivalent to the property that $\|P_{t}f\|_{L^{q(t)}(\pi)}\leq\|f\|_{L^{p}(\pi)}$ for any function $f\in L^{p}(\pi)$, where $q(t)$ is an increasing function with $q(0)=p$. For instance, the continuous-time Langevin diffusion satisfies this property, but it is not immediately clear whether such hypercontractivity is preserved upon discretization. Recent advancements, such as [CEL+24], showed that discretized Langevin processes retain hypercontractivity under LSI when the target $\pi$ is positive and log-smooth.
Nonetheless, extending hypercontractivity algorithmically beyond this favorable setting remains nontrivial. In particular, circumstances involving constraints—such as sampling from a convex body—or more general smoothness assumptions (e.g., $B_{1}\subseteq\mathcal{K}$) have resisted straightforward solutions. One might consider using reflected Langevin diffusions, but discretizations like the projected Langevin algorithm [BEL18] require access to a projection oracle for $\mathcal{K}$, and existing analytical guarantees similarly collapse to $\mathsf{TV}$-distance. Thus, achieving hypercontractivity in broader contexts remains as an open and intriguing direction.
Hypercontractivity under simultaneous heat flow (§3).
To address the boosting question, we explore an alternative strategy involving a different Markov semigroup that can be algorithmically realized via $\mathsf{PS}$. Specifically, we aim to establish hypercontractivity for the $L^{2}(\nu)$-adjoint $Q_{t}$ of the heat semigroups $P_{t}$, under ($\mathsf{LSI}$) for the target measure $\nu$. Formally, we state the following hypercontractivity result for $Q_{t}$:
@@
Let $\nu$ be a probability measure satisfying ($\mathsf{LSI}$) with $C_{\mathsf{LSI}}(\nu)<\infty$. We denote $\nu_{t}=\nu*\gamma_{t}$ for $t\geq 0$, and define $q:[0,\infty)\to(1,\infty)$ by
@@
Then, the $L^{2}(\nu)$-adjoint $Q_{t}$ of heat semigroups satisfies
@@
Its immediate consequence follows by selecting $f=\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$ for a probability measure $\mu\ll\nu$. Since $Q_{t}f=\frac{P_{t}(f\nu)}{P_{t}\nu}$, this yields $\lVert\mathrm{d}\mu_{t}/\mathrm{d}\nu_{t}\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert\mathrm{d}\mu/\mathrm{d}\nu\rVert_{L^{p}(\nu)}$, which implies $\mathcal{R}_{q(t)}(\mu_{t}\mathbin{\|}\nu_{t})\leq\mathcal{R}_{p}(\mu\mathbin{\|}\nu).$ This result has algorithmic implications for $\mathsf{PS}$. Below, the constant $2$ can be replaced by any number larger than $1$.
Corollary 1.3 (Hypercontractivity of proximal sampler).
Let $\pi_{k}^{X}$ be the law of the $k$-th iterate returned by the $\mathsf{Proximal\ sampler}$ with step size $h>0$, initial $\pi_{0}^{X}$, and target $\pi^{X}$. Given $1<p<q$, we can ensure $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\mathcal{R}_{p}(\pi_{0}^{X}\mathbin{\|}\pi^{X})$ after $N\lesssim h^{-1}C_{\mathsf{LSI}}(\pi^{X})\log_{2}\frac{q-1}{p-1}$ iterations. Moreover, given $\varepsilon>0$ and $q_{0},q>1$, we can attain $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
This result translates hypercontractivity into practical algorithmic outcomes, enabling efficient implementation of warm-start generation. Moreover, it leads to $\widetilde{O}(\kappa d\log\frac{q}{\varepsilon})$-mixing of $\mathsf{PS}$ in the well-conditioned setting (see Proposition 3.2), improving the linear dependence of $\mathsf{PS}$ on $q$ [CCSW22].
Streamlined order-preserving warm-start generation (§4).
We streamline the design and analysis of a known Gaussian-annealing scheme for warm-start generation, incorporating the hypercontractivity established for $\mathsf{PS}$. This streamlined approach offers an elegant and clean picture for warm-start generation via annealing schemes.
As with $\mathsf{PS}_{\textup{unif}}$, we propose and analyze its Gaussian variant in §4.1, denoted by $\mathsf{PS}_{\textup{Gauss}}$, specifically tailored for truncated Gaussian distributions. Leveraging the hypercontractivity result for $\mathsf{PS}$ and the bound $C_{\mathsf{LSI}}(\gamma_{\sigma^{2}}|_{\mathcal{K}})\leq\sigma^{2}$, we derive an enhanced guarantee of $\mathsf{PS}_{\textup{Gauss}}$.
Theorem 1.4.
Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body with $B_{1}(0)\subset\mathcal{K}$, given access to the membership oracle. For $q_{0}>1$, let $\mu_{k}$ denote the law of the $k$-th iterate of $\mathsf{PS}_{\textup{Gauss}}$, initialized from $\mu_{0}$ toward the target $\mu=\gamma_{\sigma^{2}}|_{\mathcal{K}}$, with step size $h\asymp(d^{2}\,\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu))^{-1}$ and threshold $\tau\asymp\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)\,e^{\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}$. Then, for any $\varepsilon>0$ and $q>1$, we can ensure $\mathcal{R}_{q}(\mu_{N}\mathbin{\|}\mu)\leq\varepsilon$ after $N\asymp d^{2}\sigma^{2}\log\frac{q\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}{(2\wedge q_{0}-1)\,\varepsilon}$ iterations, using
$2\tau N=\widetilde{O}\bigl{(}d^{2}\sigma^{2}\,e^{\mathcal{R}_{2\wedge q_{0}}(\mu_{0}\mathbin{\|}\mu)}\log\frac{q}{(2\wedge q_{0}-1)\,\varepsilon}\bigr{)}\qquad\text{membership queries in expectation}\,.$
Equipped with $\mathsf{PS}_{\textup{Gauss}}$, our warm-start generation process is very simple: (1) sample from $\mu_{1}\propto\pi_{\mathcal{K}}\gamma_{d^{-1}}$ by rejection sampling with proposal $\gamma_{d^{-1}}$, and (2) while $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ toward $\mu_{i+1}\propto\pi_{\mathcal{K}}\gamma_{\sigma_{i+1}^{2}}$ using $\mathsf{PS}_{\textup{Gauss}}$, where $\mathcal{R}_{3}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$ and $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,(1+\nicefrac{{\sigma_{i}}}{{R}})$.
This simpler algorithm eliminates unnecessary pre-processing steps, such as truncation of the convex body, and post-processing, such as adjusting target accuracy. Moreover, a key practical advantage of our streamlined method is the introduction of early-stopping at $\sigma^{2}\approx qR\Lambda^{1/2}\log^{1/2}d$. Consequently, faster-sampling phases beyond $\sigma^{2}\gtrsim R\Lambda^{1/2}\log^{1/2}d$ in [KV25a] become redundant.
On theoretical side, hypercontractivity enables the propagation of $\mathcal{R}_{q}$-warmness throughout the annealing scheme. Furthermore, we demonstrate in §4.2 that closeness between successive annealing distributions follows directly from Rényi versions of ($\mathsf{LSI}$) and ($\mathsf{PI}$), leading to the following comprehensive result for warm-start generation:
Theorem 1.5.
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ with $B_{1}(0)\subset\mathcal{K}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$, given access to the membership oracle for $\mathcal{K}$. For any $q\geq 2$, there exists an annealing algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)=O(1)$, using $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})$ membership queries in total.
1.1.3 Log-concave sampling under an evaluation oracle
Algorithm Type Complexity Reference
[…]
$\mathsf{Ball\ walk}$ $\mathcal{R}_{\infty}\to\mathsf{TV}$ $\frac{M_{\infty}^{4}d^{2}R^{2}}{\varepsilon^{4}}$ [LV07]
$\mathsf{Hit\text{-}and\text{-}Run}$ $\mathcal{R}_{\infty}\to\mathsf{TV}$ $\frac{M_{\infty}^{4}d^{2}R^{2}}{\varepsilon^{4}}$ [LV07]
$\mathcal{R}_{2}\to\mathcal{R}_{2}$ $d^{2}R^{2}\operatorname{polylog}\frac{M_{2}}{\varepsilon}$ [LV06a]
$\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ $\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ $qM_{\infty}d^{2}(\Lambda\vee 1)\operatorname{polylog}\frac{1}{\eta\varepsilon}$ [KV25b]
$\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ $M_{\infty}d^{2}R^{2}\operatorname{polylog}\frac{1}{\eta\varepsilon}$
$\mathcal{R}_{\widetilde{O}(1)}\to\mathcal{R}_{2}$ $M_{\widetilde{O}(1)}d^{2}(\Lambda\vee 1)\operatorname{polylog}\frac{1}{\eta\varepsilon}$ [KV25a]
$\mathsf{PS}_{\textup{exp}}$ $\mathcal{R}_{q}\to\mathcal{R}_{q}$ $qM_{q}^{q/(q-1)}d^{2}(\Lambda\vee 1)\log\frac{1}{\varepsilon}$ Theorem 1.6
[…]
The uniform-sampling results extend naturally to general log-concave distributions $\pi\propto e^{-V}$, given access to an evaluation oracle for a convex function $V:\mathbb{R}^{d}\to\mathbb{R}\cup\{\infty\}$. In this setting, the regularity condition $B_{1}\subset\mathcal{K}$ translates to a condition on a level set of $V$, namely, that the ground set $\mathsf{L}_{\pi,g}:=\{x\in\mathbb{R}^{d}:V(x)-\min V\leq 10d\}$ contains a unit ball. This generalization is justified by the fact that $\mathsf{L}_{\pi,g}$ takes up a constant fraction of the $\pi$-measure (see [KV25b, Lemma 2.6]).
[…]
Recent work [KV25b, KV25a] revisited this problem by adapting $\mathsf{PS}$ to the general log-concave setting. Although $\mathsf{PS}$ can in principle be extended to any distribution where $\pi^{X|Y=y}\propto\pi^{X}(\cdot)\,\gamma_{h}(\cdot-y)$ is readily samplable (e.g., when given access to a proximal oracle in the log-smooth setting), it is unclear how to implement this step for arbitrary $\pi^{X}$. To address this, [KV25b] introduced the exponential lifting technique: for $\pi^{X}\propto e^{-V}$, consider the lifted distribution
Here, the $X$-marginal of $\pi^{X,T}$ is exactly $\pi^{X}$, and $\mathcal{K}$ is convex whenever $\pi^{X}$ is log-concave. Thus, log-concave sampling reduces to the simpler problem of sampling from a log-concave distribution with a linear potential over a convex set. We then apply $\mathsf{PS}$ to this exponential distribution, denoted by $\mathsf{PS}_{\textup{exp}}$: for $z=(x,t)\in\mathbb{R}^{d+1}$ and $y\in\mathbb{R}^{d+1}$, it alternates $y\sim\pi^{Y|Z=z}=\mathcal{N}(z,hI_{d+1})$ and $z\sim\pi^{Z|Y=y}\propto\mathcal{N}(y-h\alpha,hI_{d+1})|_{\mathcal{K}}$ for $\alpha=de_{d+1}$, where the second step is implemented via rejection sampling, using the proposal $\mathcal{N}(y-h\alpha,hI_{d+1})$ with threshold $\tau$.
In [KV25b], it was shown that $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{\infty}\to\mathcal{R}_{q}$ with query complexity $\widetilde{O}(qM_{\infty}d^{2}(\Lambda\vee 1)\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$ and $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{\infty}\to\mathcal{R}_{\infty}$ with complexity $\widetilde{O}(M_{\infty}d^{2}(d\vee R^{2})\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$. As in the uniform sampling case, the order of initial warmness was relaxed to $\mathcal{R}_{c}$ with $c=\widetilde{O}(1)$ in [KV25a], where $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{c}\to\mathcal{R}_{2}$ achieves query complexity $\widetilde{O}(M_{c}d^{2}(\Lambda\vee 1)\operatorname{polylog}\nicefrac{{1}}{{\eta\varepsilon}})$. However, they were unable to balance the orders of initial warmness and final guarantee due to intricate interdependencies among algorithmic parameters.
Extension to log-concave sampling (§5).
Our redesigned $\mathsf{PS}_{\textup{exp}}:\mathcal{R}_{q}\to\mathcal{R}_{q}$ resolves this issue, aligning the order of initial warmness with that of the final guarantee. At an additive cost of $qd^{2}$ queries, our result effectively tames a long line of work on sampling in this classical setting. See §5.1 for further details.
Theorem 1.6 (Zeroth-order log-concave sampling from a warm start).
Let $\pi\propto e^{-V}$ be a log-concave distribution over $\mathbb{R}^{d}$ with $B_{1}\subset\mathsf{L}_{\pi,g}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and an evaluation oracle for $V$. Let $\pi_{k}$ be the $X$-marginal of the $k$-th iterate of $\mathsf{PS}_{\textup{exp}}$ with initial distribution $\pi_{0}$, step size $h\asymp(d^{2}\,\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi))^{-1}$, and threshold $\tau\asymp\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)\,e^{\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)}$ for $q\geq 2$. Given $\varepsilon>0$, we can ensure $\mathcal{R}_{q}(\pi_{N}\mathbin{\|}\pi)\leq\varepsilon$ with $N\asymp qd^{2}(\Lambda\vee 1)\,\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)\log d\log\frac{1}{\varepsilon}$, using
$2\tau N=\widetilde{O}\bigl{(}qd^{2}(\Lambda\vee 1)\,e^{\mathcal{R}_{q}(\pi_{0}\mathbin{\|}\pi)}\log\frac{1}{\varepsilon}\bigr{)}\qquad\text{evaluation queries in expectation}\,.$
We remark on a connection to the well-conditioned setting. In this case, the ground set contains a ball of radius $\Theta((d/\beta)^{1/2})$ centered at $\arg\min V$. Since $\Lambda\leq\alpha^{-1}$ by the Brascamp–Lieb or Lichnerowicz inequality, $\mathsf{PS}_{\textup{exp}}$ uses $\widetilde{O}(q\,(\kappa d+d^{2})\log\nicefrac{{1}}{{\varepsilon}})$ queries from an $O(1)$-warm start in $\mathcal{R}_{q}$. With the additional $qd^{2}$ queries, this matches the known $\widetilde{O}(q\kappa d\log\nicefrac{{1}}{{\varepsilon}})$-complexity results for zeroth-order samplers in this setting when additionally given a proximal oracle. See §1.3 for further details.
Annealing approaches for uniform sampling have also been extended to arbitrary log-concave distributions [LV06a, KV25b, KV25a]. In the general case, the best-known complexity for warm-start generation remains comparable to that of the convex-body case for $\mathsf{TV}$ and $\mathcal{R}_{\infty}$-warmness, with algorithms requiring additional $d^{2}\operatorname{polylog}\nicefrac{{1}}{{\varepsilon}}$ queries. In §5.2, we extend our order-preserving warm-start generation to general log-concave distributions in a similar fashion:
Theorem 1.7.
Let $\pi\propto e^{-V}$ be a log-concave distribution over $\mathbb{R}^{d}$ with $B_{1}(0)\subset\mathsf{L}_{\pi,g}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, $\mathbb{E}_{\pi}[|\cdot|^{2}]=R^{2}$, and access to an evaluation oracle for $V$. For $q\geq 2$, there exists an annealing algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)=O(1)$, using $\widetilde{O}(d^{2.5}+qd^{2}(R^{3/2}\vee 1)(\Lambda^{1/4}\vee 1))$ evaluation queries in total.
1.2 Technical overview
Proximal sampler with restart.
Compared to the standard $\mathsf{PS}$, our approach has a key modification—restarting the current iteration if the number of attempts in rejection sampling for the backward step exceeds a fixed threshold. The analysis of $\mathsf{PS}$ is split to two parts: the number of iterations required to achieve desired sampling accuracy, and the per-iteration query complexity arising from rejection sampling. Both components have been well-studied in prior works. The convergence rate of $\mathsf{PS}$ under ($\mathsf{PI}$) was established in [CCSW22], along with a rigorous justification for the general de Bruijn identity [KO25]. Moreover, [KVZ24] proposed carefully tuned parameter choices for $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$ that keep the per-step complexity under control.
Given these foundations, the main technical questions concern the effect of the restarting mechanism. We must ensure that restarting neither alters the transition kernel of $\mathsf{PS}$ nor introduces excessive overhead in query complexity. The first point is intuitively straightforward: each iteration of $\mathsf{PS}$ can be viewed as sampling from the true transition kernel followed by an acceptance step. The distribution of the first successful sample thus corresponds to the intended transition kernel (Lemma 2.7). The second point follows from the same perspective. Suppose that each iteration succeeds with probability $p$, and the threshold for rejection retries is set to $\tau$. Then, the expected number of retries until a successful iteration is bounded by $\tau/p$. This bound remains moderate so long as the success probability is bounded away from zero, allowing us to employ simpler choices of step size $h$ and threshold $\tau$ than those used in $\mathsf{\mathsf{In\text{-}and\text{-}Out}}$.
Hypercontractivity under heat flow.
@@ 1.2 Technical overview
In our analysis of Gaussian annealing (§4.2), we present a more principled and streamlined approach to bounding the closeness between successive annealing distributions. Whereas prior works relied on ad-hoc arguments to establish such closeness, we show that this property follows more cleanly from LSI, combined with a mollification argument that allows for a limiting process.

### v2

@@ 1.1.2 Improved warm-start generation via annealing
Looking back at diffusion for inspiration, perhaps the most relevant mathematical notion to tackle this question is Gross’ hypercontractivity of semigroups [GRO75] (see Proposition 3.1): for a semigroup $(P_{t})_{t\geq 0}$ reversible with respect to an invariant measure $\pi$, it is known that ($\mathsf{LSI}$) for $\pi$ is equivalent to the property that $\|P_{t}f\|_{L^{q(t)}(\pi)}\leq\|f\|_{L^{p}(\pi)}$ for any function $f\in L^{p}(\pi)$, where $q(t)$ is an increasing function with $q(0)=p$. For instance, the continuous-time Langevin diffusion satisfies this property, but it is not immediately clear whether such hypercontractivity is preserved upon discretization. Recent advancements, such as [CEL+24], showed that discretized Langevin processes retain hypercontractivity under LSI when the target $\pi$ is unconstrained and log-smooth.
Can we extend hypercontractivity algorithmically beyond such specialized settings? In particular, sampling with hard constraints—such as sampling from a convex body—or more general smoothness assumptions (e.g., $B_{1}\subseteq\mathcal{K}$) have resisted straightforward solutions. One might consider using reflected Langevin diffusions, but discretizations like the projected Langevin algorithm [BEL18] require access to a projection oracle for $\mathcal{K}$, and existing analytical guarantees similarly collapse to $\mathsf{TV}$-distance (beside incurring much higher complexity).
Hypercontractivity of the heat adjoint (§3).
To address this boosting question, we explore an alternative strategy involving a different Markov semigroup that can be algorithmically realized via $\mathsf{PS}$. Specifically, we aim to establish hypercontractivity for the $L^{2}(\nu)$-adjoint $Q_{t}$ of the heat semigroup $P_{t}$, under ($\mathsf{LSI}$) for the target measure $\nu$. Formally, we state the following hypercontractivity result for $Q_{t}$.
@@
Let $\nu$ be a probability measure satisfying ($\mathsf{LSI}$) with $C_{\mathsf{LSI}}(\nu)<\infty$. Let $\nu_{t}=\nu*\gamma_{t}$ for $t\geq 0$, and define $q:[0,\infty)\to(1,\infty)$ by
@@
Then, the $L^{2}(\nu)$-adjoint $Q_{t}$ of the heat semigroup satisfies
@@
(1) Proof sketch.
@@
(2) Consequences.
An immediate consequence follows by selecting $f=\frac{\mathrm{d}\mu}{\mathrm{d}\nu}$ for a probability measure $\mu\ll\nu$. Since $Q_{t}f=\frac{P_{t}(f\nu)}{P_{t}\nu}$, this yields $\lVert\mathrm{d}\mu_{t}/\mathrm{d}\nu_{t}\rVert_{L^{q(t)}(\nu_{t})}\leq\lVert\mathrm{d}\mu/\mathrm{d}\nu\rVert_{L^{p}(\nu)}$, which implies $\mathcal{R}_{q(t)}(\mu_{t}\mathbin{\|}\nu_{t})\leq\mathcal{R}_{p}(\mu\mathbin{\|}\nu).$ This result has algorithmic implications for $\mathsf{PS}$. Below, the constant $2$ can be replaced by any number larger than $1$.
Corollary 1.3 (Hypercontractivity of proximal sampler).
Let $\pi_{k}^{X}$ be the law of the $k$-th iterate returned by the $\mathsf{Proximal\ sampler}$ with step size $h>0$, initial $\pi_{0}^{X}$, and target $\pi^{X}$. Given $1<p\leq q$, we have $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\mathcal{R}_{p}(\pi_{0}^{X}\mathbin{\|}\pi^{X})$ after $N\lesssim h^{-1}C_{\mathsf{LSI}}(\pi^{X})\log_{2}\frac{q-1}{p-1}$ iterations. Moreover, given $\varepsilon>0$ and $q\geq q_{0}>1$, we have $\mathcal{R}_{q}(\pi_{N}^{X}\mathbin{\|}\pi^{X})\leq\varepsilon$ after
This result translates hypercontractivity into an algorithmic outcome, enabling efficient implementation of warm-start generation. Moreover, it leads to $\widetilde{O}(\kappa d\log\frac{q}{\varepsilon})$-mixing of $\mathsf{PS}$ in the well-conditioned setting (Proposition 3.2), improving the linear dependence of $\mathsf{PS}$ on $q$ [CCS+22].
Streamlined warm-start generation (§4).
We simplify the design and analysis of a known Gaussian-annealing scheme for warm-start generation, incorporating the hypercontractivity established for $\mathsf{PS}$. This streamlined approach offers an elegant and clean picture for warm-start generation via annealing schemes.
As with $\mathsf{PS}_{\textup{unif}}$, we propose and analyze its Gaussian variant in §4.1, denoted by $\mathsf{PS}_{\textup{Gauss}}$, specifically tailored for truncated Gaussian distributions $\gamma_{\sigma^{2}}|_{\mathcal{K}}$. Leveraging the hypercontractivity result for $\mathsf{PS}$ and the bound $C_{\mathsf{LSI}}(\gamma_{\sigma^{2}}|_{\mathcal{K}})\leq\sigma^{2}$, we derive an enhanced guarantee for $\mathsf{PS}_{\textup{Gauss}}$. Below, $B_{1}(0)$ denotes the $\ell_{2}$ unit ball centered at the origin.
Theorem 1.4.
Let $\mathcal{K}\subset\mathbb{R}^{d}$ be a convex body with $B_{1}(0)\subset\mathcal{K}$ given by a membership oracle, and let $\mu\propto\gamma_{\sigma^{2}}\cdot\mathds{1}_{\mathcal{K}}$ be the Gaussian truncated to $\mathcal{K}$. Given $\varepsilon\in(0,1/100)$, initial distribution $\mu_{0}$, any $q\geq 2\vee\widetilde{\Omega}(\log(d^{2}\sigma^{2}\log\frac{1}{\varepsilon}))$ such that $q\geq q_{0}:=2\vee\widetilde{O}(\log(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log\frac{1}{\varepsilon}))$ with $M_{q_{0}}=\lVert\frac{\mathrm{d}\mu_{0}}{\mathrm{d}\mu}\rVert_{L^{q_{0}}(\mu)}\leq 10$, consider the algorithm that runs $\mathsf{PS}_{\textup{Gauss}}$ initialized at $\mu_{0}$ and restarts if $\mathsf{PS}_{\textup{Gauss}}$ fails before completing $N=\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon})$ iterations. There exist choices of $h$ and $\tau$ such that this algorithm restarts with probability at most $1/50$, and its output $Z$ satisfies $\mathcal{R}_{q}(\operatorname{law}Z\mathbin{\|}\mu)\leq\varepsilon$, using $\widetilde{O}(d^{2}\sigma^{2}\log\frac{q}{\varepsilon}\log^{5}\frac{1}{\varepsilon})$ membership queries in expectation.
Equipped with $\mathsf{PS}_{\textup{Gauss}}$, our warm-start generation process is very simple: (1) sample from $\mu_{1}\propto\pi_{\mathcal{K}}\gamma_{d^{-1}}$ by rejection sampling with the Gaussian proposal $\gamma_{d^{-1}}$, and (2) while $\sigma_{i}^{2}\lesssim qR\Lambda^{1/2}\log^{1/2}d$, move from $\bar{\mu}_{i}$ toward $\mu_{i+1}\propto\pi_{\mathcal{K}}\gamma_{\sigma_{i+1}^{2}}$ using $\mathsf{PS}_{\textup{Gauss}}$, where $\mathcal{R}_{\widetilde{O}(1)}(\bar{\mu}_{i}\mathbin{\|}\mu_{i})=O(1)$ and $\sigma_{i+1}^{2}=\sigma_{i}^{2}\,(1+\nicefrac{{\sigma_{i}}}{{R}})$.
This simpler algorithm eliminates pre-processing steps, such as truncation of the convex body, and post-processing, such as adjusting the target accuracy. Moreover, a key practical advantage of this streamlined method is the introduction of early-stopping at $\sigma^{2}\approx qR\Lambda^{1/2}\log^{1/2}d$. Consequently, faster sampling phases beyond $\sigma^{2}\gtrsim R\Lambda^{1/2}\log^{1/2}d$ needed in [KV25a] become redundant. On the theoretical side, hypercontractivity enables the propagation of $\mathcal{R}_{q}$-warmness throughout the annealing scheme, leading to the following comprehensive result for warm-start generation:
Theorem 1.5.
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ given by a membership oracle with $B_{1}(0)\subset\mathcal{K}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$, . For any $\varepsilon_{\textup{unif}}\in(0,1/100)$ and $q\geq 2\vee\widetilde{\Omega}\bigl(\log\frac{dR}{\varepsilon_{\textup{unif}}}\bigr)$, there exists an algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)\leq 2$, using $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4})$ membership queries in expectation.
Combining this with Theorem 1.1, we obtain the query complexity of convex-body sampling from scratch, with $\mathcal{R}_{q}$-divergence output guarantees for any $q\geq 1$.
Corollary 1.6.
Let $\pi$ be the uniform distribution over a convex body $\mathcal{K}\subset\mathbb{R}^{d}$ given by a membership oracle with $B_{1}(0)\subset\mathcal{K}$, $\Lambda=\lVert\operatorname{cov}\pi\rVert$, and $R^{2}=\mathbb{E}_{\pi}[|\cdot|^{2}]$. For any $\varepsilon\in(0,1/100)$ and $q\geq 1$, there exists an algorithm that outputs a sample $X^{*}$ such that $\mathcal{R}_{q}(\operatorname{law}X^{*}\mathbin{\|}\pi)\leq\varepsilon$, using $\widetilde{O}(qd^{2}R^{3/2}\Lambda^{1/4}\log\frac{1}{\varepsilon}+qd^{2}\Lambda\log^{7}\frac{1}{\varepsilon})$ membership queries in expectation.
@@
In our analysis of Gaussian annealing (§4.2), we present a principled and streamlined approach to bounding the closeness between successive annealing distributions. Whereas prior works relied on localization arguments (with long calculations) to establish such closeness, we show that this property follows more cleanly from LSI, combined with a mollification argument that allows for a limiting process.